OpenStudy (anonymous):

Continuous random variable X has density given by f(x)=c⋅(x+1)*(x−3) if x∈[−1,3] and 0 otherwise. 1) Find DF(2;X) . distribution function 2) Find E[X] .

OpenStudy (zarkon):

what have you tried?

OpenStudy (anonymous):

i try to use the integral by putting this together as x^2-2x-3

OpenStudy (irishboy123):

\(f(x)=c(x+1)(x−3)\) \(F(x) = \int\limits_{-1}^{3} \; dx \quad c(x^2-2x-3)\)

OpenStudy (anonymous):

\[\frac{ x^3 }{ 3 }- x^2 -3x\]

OpenStudy (irishboy123):

for the next bit, you might try \(G(x) = \int\limits_{-1}^{3} \; x \; dx \quad c(x^2-2x-3) = \int\limits_{-1}^{3} \; dx \quad c(x^3-2x^2-3x)\)

OpenStudy (anonymous):

ok

OpenStudy (zarkon):

you need to find the value of \(c\) such that \[\int\limits_{-1}^{3} c(x^2-2x-3)dx=1\]

OpenStudy (anonymous):

is